28- Auerbach Inc. issued 10% bonds on October 1, 2021. The bonds have a maturity date of September 30, 2031 and a face value of $300 million. The bonds pay interest each March 31 and September 30, beginning March 31, 2022. The effective interest rate established by the market was 12%.

Assuming that Auerbach issued the bonds for $265,590,000, what would the company report for its net bond liability balance at December 31, 2021, rounded up to the nearest thousand? (Do not round intermediate calculations.)

Multiple Choice

  • $273,090,000.

  • $258,090,000.

  • $299,000,000.

  • $266,058,000.
Answer #1

Issue price of bonds = 265590000

Face value of bonds = 300000000

Interest expense = 265590000*12%*3/12 = 7967700

Interest paid = 300000000*10%*3/12 = 7500000

Amortization = 7967700-7500000 = 467700

Net bond liability = 265590000+467700 = 266058000

So answer is d) $266058000